Operations Report – May 11th

Crew 212 Operations Report 11-MAY-2019

SOL: 6

Name of person filing report: Paolo Guardabasso

Non-nominal systems: None

Notes on non-nominal systems: Nothing to report

Generator (hours run last night): 9:30

On (last night): 9:00 PM

Off (this morning): 6:30 AM

Solar— SOC 88% at 7:30 PM

Diesel Reading – 95% [From Director]

Propane Reading – 45%

Ethanol Free Gasoline – 4 gallons

Water (Axillary tank) – 0 gallons

Water (static tank) – 170 gallons

Axillary to Static tank transfer – no

Gallons transferred: No

Water in GreenHab – 150 gallons

Water (loft) – Static to Loft Pump used – Yes

Water Meter: 143500.4

Toilet tank emptied: Yes

Deimos rover used: no (out of order)


Beginning charge:

Ending charge:

Currently charging:

Sojourner rover used: ASSIGNED TO DIRECTOR


Beginning charge:

Ending charge:

Currently charging:

Spirit rover used: Yes

Hours: 92.8h

Beginning charge: 100%

Ending charge: 79%

Currently charging: Yes

Opportunity rover used: Yes

Hours: 59.1h

Beginning charge:

Ending charge:

Currently charging: Yes

Curiosity rover used: Yes


Beginning charge:

Ending charge:

Currently charging: Yes

Notes on rovers: Curiosity and Opportunity were missing one wheel hubcap per rover. It’s not yet clear exactly how they went missing but the Director is dealing with the matter and the one on Curi has already been replaced. A new charging cable was added in order to be able to charge 3 rovers at the same time.

ATV’s Used: (Honda, 300, 350.1, 350.2, 350.3)

Reason for use:

Oil Added?

ATV Fuel Used: Gals

# Hours the ATVs were Used today:

Notes on ATVs:

HabCar used and why, where? Nothing to report

CrewCar used and why, where? Off-Campus

General notes and comments: The crew worked on their experiment and an EVA was delayed and lasted 1 hour instead of 2.

Summary of internet: Nothing to report.

Summary of suits and radios: Some radios didn’t hold their charge, major testing will be carried out in the coming days. Radios proved not working at a range of about 1 km from the Hab during EVA. Some spare batteries for suits were not able to recharge.

Summary of Hab operations: Fixing of the oven door is still ongoing.

Summary of GreenHab operations: The fan was turned on as the temperature passed 40° C and it managed to cool the room to 30°C.

Summary of ScienceDome operations: Microbiological experiments were run.

Summary of RAM operations: Nothing to report.

Summary of any observatory issues: The crew astronomer reported that the telescope is not properly focused.

Summary of health and safety issues: Nothing to report.

Questions, concerns, and requests to Mission Support: Several tourist cars approached the MDRS; the one that passed the sign was stopped by the Director on its way to the Hab.

Copyright © The Mars Society. All rights reserved. | Main Site